OpenStudy (anonymous):

x^4 - 7x^2 -18 = 0 Solve

OpenStudy (anonymous):

x4−7x2−18=0 Step 1: Factor left side of equation. (x2+2)(x+3)(x−3)=0 Step 2: Set factors equal to 0. x2+2=0 or x+3=0 or x−3=0 x=−3 or x=3